# Tuning-Free Alignment of Diffusion Models with Direct Noise Optimization

### v1

@@ 6 Conclusions
In this work, we present a comprehensive study on Direct Noise Optimization (DNO) for aligning diffusion generative models. We introduce variants of DNO designed to efficiently address challenges such as out-of-distribution reward-hacking and the optimization of non-differentiable reward functions. More significantly, we demonstrate the exceptional efficacy of DNO, underscoring its capacity to rival tuning-based methods. The primary limitation of DNO lies in its integration with the sampling process of diffusion models, leading to a substantial increase in processing time compared to direct sampling. Nonetheless, we argue that the additional time cost, being within a reasonable and acceptable range, is a worthwhile trade-off for attaining high-reward generated samples in a wide range of real-world applications. We anticipate DNO gaining greater attention in future research and applications due to its flexibility to accommodate any reward function—or even a combination of different reward functions—while demanding only modest computing resources, positions it as an accessible tool for many practitioners.
@@ C Empirical Investigation of $P(z)$
Firstly, in Figure 13(a), we revisit the examples from Figures 4(a) and 4(b), coloring each sample based on the value of $P(z)$. As depicted in Figure 13(a), $P(z)$ proves to be an efficient metric to separate in-distribution samples from out-of-distribution samples; those in-distribution have high values for $P(z)$, whereas those out-of-distribution exhibit values of $P(z)$ near zero.
@@
We further visualize the entire optimization trajectory for the examples in Figures 4(c) and 4(d), i.e., optimizing the whiteness reward for Stable Diffusion v1.5 with the prompt "black duck" in Figure 14. Specifically, from Figure 14 we can clearly see that the value of $P(z)$ gradually decreases, and the generated image also gradually diverges from the distribution associated with a black duck. Notably, at around $20$ steps, the value of $P(z)$ becomes near-zero, and at the same time, the generated image more closely resembles a blue duck rather than the specified black duck.
@@ D.1 Algorithm Implementation
It is clear that to solve the direct noise optimization problem stated in Problem 7, differentiation of the noise-to-sample mapping $M_{\theta}$ is required. It is worth noting that this differentiation cannot be handled by standard auto-differentiation in PyTorch [Paszke et al., 2019], as it can lead to a memory explosion. A common technique to resolve this issue is gradient checkpointing, which has also been adopted by other related works on noise optimization [Wallace et al., 2023b, Novack et al., 2024, Karunratanakul et al., 2023].

### v2

@@ 6 Conclusions
In this work, we present a comprehensive study on Direct Noise Optimization (DNO) for aligning diffusion generative models. We introduce variants of DNO designed to efficiently address challenges such as out-of-distribution reward-hacking and the optimization of non-differentiable reward functions. More significantly, we demonstrate the exceptional efficacy of DNO, underscoring its capacity to rival tuning-based methods. The primary limitation of DNO lies in its integration with the sampling process of diffusion models, leading to a substantial increase in processing time compared to direct sampling. Nonetheless, we argue that the additional time cost, being within a reasonable and acceptable range, is a worthwhile trade-off for attaining high-reward generated samples in a wide range of real-world applications. We anticipate DNO gaining greater attention in future research and applications due to its flexibility to accommodate any reward function—or even a combination of different reward functions—while demanding only modest computing resources, which positions it as an accessible tool for many practitioners.
@@ C Empirical Investigation of $P(z)$
Firstly, in Figure 15(a), we revisit the examples from Figures 6(a) and 6(b), coloring each sample based on the value of $P(z)$. As depicted in Figure 15(a), $P(z)$ proves to be an efficient metric to separate in-distribution samples from out-of-distribution samples; those in-distribution have high values for $P(z)$, whereas those out-of-distribution exhibit values of $P(z)$ near zero.
@@
We further visualize the entire optimization trajectory for the examples in Figures 6(c) and 6(d), i.e., optimizing the whiteness reward for Stable Diffusion v1.5 with the prompt "black duck" in Figure 16. Specifically, from Figure 16 we can clearly see that the value of $P(z)$ gradually decreases, and the generated image also gradually diverges from the distribution associated with a black duck. Notably, at around $20$ steps, the value of $P(z)$ becomes near-zero, and at the same time, the generated image more closely resembles a blue duck rather than the specified black duck.
@@ D.1 Algorithm Implementation
It is clear that to solve the direct noise optimization problem stated in Problem 8, differentiation of the noise-to-sample mapping $M_{\theta}$ is required. It is worth noting that this differentiation cannot be handled by standard auto-differentiation in PyTorch [Paszke et al., 2019], as it can lead to a memory explosion. A common technique to resolve this issue is gradient checkpointing, which has also been adopted by other related works on noise optimization [Wallace et al., 2023b, Novack et al., 2024, Karunratanakul et al., 2023].